OpenStudy (nsh4267):

Solve the equation log(x^4) = (log x)^2

OpenStudy (anonymous):

\[\log(x^4) = (\log(x))^2\]

OpenStudy (anonymous):

Power rule: \(\log(x^a) = a \cdot \log(x)\)

OpenStudy (anonymous):

\[4 \cdot \log(x) = \log(x) \times \log(x)\]

OpenStudy (mathmath333):

\(\large\tt \begin{align} \color{black}{log(x^4) = \left( log (x)\right )^2\\~\\ 4logx = log( x)\times log (x)\\~\\ 4 = log (x)\\~\\ log (x)=4 \\~\\ x=10^4 \\~\\ x=10000 \\~\\ }\end{align}\)

OpenStudy (nsh4267):

It still says it is wrong

OpenStudy (anonymous):

really?

OpenStudy (anonymous):

What is the base??

OpenStudy (mathmath333):

it might be natural; log which is represented as \(\Large ln(x)\) \(\large\tt \begin{align} \color{black}{ln(x^4) = (ln x)^2\\~\\ 4lnx = ln x\times ln x\\~\\ 4 = ln x\\~\\ ln x=4 \\~\\ x=e^4 --------(\rm e=2.718)\\~\\ x=54.59 \\~\\ }\end{align}\)